# We Can Hide More Bits: The Unused Watermarking Capacity in Theory and in Practice

### v1

@@ 1 Introduction
Content provenance, the act of identifying the origin and history of media, is increasingly important as generative tools for creating and editing text, audio and video content are becoming ever more accessible and powerful. The most robust and reliable content provenance solutions, including those in the C2PA industry standard (C2PA), combine cryptographically secure metadata, fingerprinting, and invisible watermarking. Invisible watermarking is the key area of interest in both academia and industry. This paper focuses on the image domain, as it is the most mature, despite techniques also existing for text (Kirchenbauer et al., 2023; Dathathri et al., 2024) and audio (Chen et al., 2023; San Roman et al., 2024). Image watermarking is already integrated into commercial models (e.g., Meta AI, Adobe Firefly and Google Imagen) and available for third-party use via open-source models (Zhu et al., 2018; Bui et al., 2023a), and is increasingly being considered for regulatory mandates (The White House, 2023; California State Leg., 2024; European Parliament and Council, 2024).
@@ 2 Related work
Deep learning based-watermarking. With the development of deep learning techniques for computer vision, it was only natural to extend them to image watermarking. Vukotić et al. (2018) proposed adversarially attacking a fixed image extractors, an idea later built upon by Fernandez et al. (2022) and Kishore et al. (2022). More popular and successful methods train purpose-built neural networks. Early convolutional models, such as those introduced by Mun et al. (2017) and Zhu et al. (2018), established the feasibility of CNN-based watermarking. Subsequently, architectures based on U-Net (Ronneberger et al., 2015) gained prominence, leveraging multi-resolution representations and residual connections to enable greater network depths. Recent work has also explored watermarking in the latent space of diffusion models Bui et al. (2023b). Advances in perceptual loss design have proven critical: careful loss selection and tuning improved both image quality (Bui et al., 2023a; Fernandez et al., 2024; Xu et al., 2025) and robustness (Tancik et al., 2020; Jia et al., 2021; Pan et al., 2024). Beyond robustness, methods for watermark localization and multi-message embedding have been proposed (Sander et al., 2025; Wang et al., 2024; Zhang et al., 2024), while add-on techniques such as adversarial training (Luo et al., 2020) and attention mechanisms (Zhang et al., 2020) further enhance performance. Although combining image generation with watermarking has also been studied (Fernandez et al., 2023; Wen et al., 2023; Kim et al., 2023; Hong et al., 2024; Ci et al., 2024), such generative approaches are beyond the scope of this work.
Theoretical capacity of watermarking. Most prior attempts for theoretical bounds on the watermarking capacity of images focused on mathematically amenable but detached from reality setups, often assuming Gaussian covers and perturbations. Costa (1983) and Cohen and Lapidoth (2002) showed that under the Gaussian cover and noise assumption, the fact that the decoder does not know the cover image does not affect the capacity: it is determined solely by the strength of the watermark and the channel noise. Chen and Wornell (2002) proposed quantization-based watermarking scheme with guaranteed capacity under Gaussian noise. While Moulin and O’Sullivan (2003) and Moulin and Koetter (2005) are slightly closer to practical setups, i.e., discrete values and distortions other than Gaussian noise, they can only factor robustness to perturbations with small magnitude (similarly to Somekh-Baruch and Merhav (2004)), something that does not apply to most geometric augmentations. Merhav (2005) considers geometric perturbations with large magnitude but only such that can be expressed as a permutation of the pixels. All these works take an information-theory-based approach: power-limited communication over a super-channel with a state that is known to the encoder. Maor and Merhav (2005) studied the connection with compression from an information-theoretic angle but their results cannot be used to obtain practical bounds. By contrast, we take a geometric approach in the current work which we find better suited for studying robustness to more realistic perturbations.
@@ 4.2 Our simplest bounds are achievable, yet models struggle to get near them
[…]
We can use this ${32}{\times}{32}\text{px}$ model to demonstrate that much higher capacities are indeed possible. As we do not consider robustness to any geometric or valuemetric transformations, simply tiling the lower-resolution watermarks can produce a watermark with much higher capacity at the same PSNR. Therefore, if we take the ${256}{\times}{256}\text{px}$ resolution as reference, we get $64\times$ the capacity at ${32}{\times}{32}\text{px}$. Hence, if we tile the 512 bit model which achieved 100% bit accuracy at ${41.7}\,\text{dB}$, we get a model with the impressive ${32{,}768}\,\text{bits}$ bits of capacity and still ${41.7}\,\text{dB}$ (as PSNR is resolution-independent). Therefore, we are getting effective capacity of ${32{,}768}\,\text{bits}$ by simply tiling the ${32}{\times}{32}\text{px}$ model. That is much closer to our bound of roughly ${600{,}000}\,\text{bits}$ but still only about ${0.167}\,\text{bpp}$.

### v2

@@ 1 Introduction
Content provenance, the act of identifying the origin and history of media, is increasingly important as generative tools for creating and editing text, audio and video content are becoming ever more accessible and powerful. The most robust and reliable content provenance solutions, including those in the C2PA industry standard (C2PA), combine cryptographically secure metadata, fingerprinting, and invisible watermarking. Invisible watermarking is the key area of interest in both academia and industry. This paper focuses on the image domain, as it is the most mature, despite techniques also existing for text (Kirchenbauer et al., 2023; Dathathri et al., 2024) and audio (Chen et al., 2023a; San Roman et al., 2024). Image watermarking is already integrated into commercial models (e.g., Meta AI, Adobe Firefly and Google Imagen) and available for third-party use via open-source models (Zhu et al., 2018; Bui et al., 2023a) .
@@ 2 Related work
Deep learning based-watermarking. With the development of deep learning techniques for computer vision, it was only natural to extend them to image watermarking. Vukotić et al. (2018) proposed adversarially attacking a fixed image extractors, an idea later built upon by Fernandez et al. (2022) and Kishore et al. (2022). A similar iterative approach, albeit lacking robustness, was attempted for steganography by Chen et al. (2023b). LISO operates by iteratively optimizing the embedding of information into images, but its performance degrades sharply under even mild compression. Notably, the only robustness-aware variant, LISO-JPEG, achieves merely $\sim$1 bpp at 19.7 dB PSNR, which is well below practical watermarking quality and is only robust to JPEG compression with CRF 80. This demonstrates that LISO is non-practical for real-world watermarking applications due to its weak robustness and low fidelity.
@@
More popular and successful methods train purpose-built neural networks. Early convolutional models, such as those introduced by Mun et al. (2017) and Zhu et al. (2018), established the feasibility of CNN-based watermarking. Subsequently, architectures based on U-Net (Ronneberger et al., 2015) gained prominence, leveraging multi-resolution representations and residual connections to enable greater network depths. Recent work has also explored watermarking in the latent space of diffusion models Bui et al. (2023b). Advances in perceptual loss design have proven critical: careful loss selection and tuning improved both image quality (Bui et al., 2023a; Fernandez et al., 2024; Xu et al., 2025) and robustness (Tancik et al., 2020; Jia et al., 2021; Pan et al., 2024). Beyond robustness, methods for watermark localization and multi-message embedding have been proposed (Sander et al., 2025; Wang et al., 2024; Zhang et al., 2024), while add-on techniques such as adversarial training (Luo et al., 2020) and attention mechanisms (Zhang et al., 2020) further enhance performance. Although combining image generation with watermarking has also been studied (Fernandez et al., 2023; Wen et al., 2023; Kim et al., 2023; Hong et al., 2024; Ci et al., 2024), such generative approaches are beyond the scope of this work.
Information-theoretic capacity of watermarking
A natural foundation for analyzing watermarking capacity is the seminal framework of Gel’fand and Pinsker (1980), which characterizes communication over channels with random parameters. In the watermarking context, the cover image acts as the channel state: it is known to the encoder (the hider) but not to the decoder. The Gel’fand–Pinsker theorem expresses the capacity of such channels as
$C=\max_{p(u|s),\,x(u,s)}\bigl[I(U;Y)-I(U;S)\bigr],$
where $S$ denotes the cover signal, $U$ an auxiliary variable chosen by the encoder, $X$ the watermarked image, and $Y$ the attacked image. This framework underlies nearly all classical information-theoretic analyses of watermarking.
Early works such as Costa (1983) and Cohen and Lapidoth (2002) assume Gaussian cover signals and Gaussian, memoryless attacks. Under these assumptions, the decoder’s ignorance of the cover does not reduce capacity: the achievable rate depends only on watermark power and noise power. Quantization-based analyses (Chen and Wornell, 2002) operate under similarly stylized Gaussian models.
More sophisticated approaches were introduced by Moulin and O’Sullivan (2003) and Moulin and Koetter (2005), who cast watermarking as a strategic game between hider and attacker and defined the notion of hiding capacity. These formulations characterize the fundamental tradeoff between embedding distortion $D_{1}$ and attack distortion $D_{2}$, and have inspired a long line of follow-up work (e.g., Somekh-Baruch and Merhav, 2004 and Merhav, 2005). However, these analyses remain grounded in Euclidean or Hamming distortion metrics and in pixelwise probabilistic models of the cover and the attack. Even extensions that allow more complex or adversarial behavior—such as pixel permutations—still operate within a per-pixel, additive-noise paradigm.
These classical works provide valuable theoretical insight, but they share the same foundational assumption: watermarking is modeled as a channel coding problem over an idealized Euclidean signal space, with attacks represented as noise in that space. As discussed next, this assumption diverges sharply from the realities of digital images and modern watermark attacks.
The need for a new framework for watermarking capacity
The classical Gel’fand and Pinsker (1980) framework treats watermarking as a per-pixel communication problem: attacks are modeled as additive or memoryless noise, distortions are measured in Euclidean or Hamming metrics, and images are assumed to be continuous signals with well-defined probability densities. The capacity expression
I(U;Y)I(U;S)I(U;Y)-I(U;S)
reveals the central limitation of this view: mutual information depends on the absolute probability distribution of natural images. Such distributions are unknown, representation-dependent (RGB vs. YCbCr, gamma correction, quantization, compression), and fundamentally inestimable. Because $I(U;Y)$ and $I(U;S)$ require these true underlying densities—not empirical or relative statistics—Gel’fand–Pinsker capacities cannot be instantiated for real images.
Moreover, the dominant failure modes of watermarking systems are not pixelwise perturbations but geometric transformations: cropping, resampling, rotations, perspective changes, and non-rigid warps that move pixels rather than modify their intensities. These structured distortions introduce strong spatial dependencies that fall far outside memoryless noise models and cannot be meaningfully captured by $\ell_{2}$ or related Euclidean metrics. Real images also inhabit discrete, quantized spaces such as $\{0,\ldots,255\}^{3}$, making differential-entropy–based Gaussian capacity formulas inapplicable. As a result, classical information-theoretic capacity predictions are tied to an idealized Euclidean signal model that digital images simply do not satisfy.
Because real attacks are predominantly geometric and because natural-image densities are inaccessible, classical information-theoretic models cannot yield realistic or actionable capacity estimates. They cannot answer practical questions such as:
How many bits can be reliably embedded in a $128\times 128$ image while remaining recoverable after rotations, cropping, or other structured geometric distortions?
Our goal is to develop a capacity theory grounded in the geometry of images—one that predicts how many bits can be reliably stored under realistic transformations such as affine and non-rigid warps, and that provides bounds and constructions reflecting the true failure modes of modern watermarking systems.
@@ 4.2 Our simplest bounds are achievable, yet models struggle to get near them
Lower-resolution training and tiling. Our experiments reveal that Video Seal does not exploit the additional degrees of freedom available at higher image resolutions. When trained at ${256}{\times}{256}\text{px}$, the model achieves essentially the same capacity and PSNR as when trained at ${32}{\times}{32}\text{px}$ (see Section˜4.1 and Table˜2). To verify this, we train Video Seal in the setup of Section˜4.1 at ${32}{\times}{32}\text{px}$ using the same learning-rate and $\lambda_{i}$ sweeps for 600 epochs. As shown in Figure˜7 (centre) and Table˜2, the performance at ${32}{\times}{32}\text{px}$ is nearly identical to that at ${256}{\times}{256}\text{px}$: the 512-bit model reaches 100% bit accuracy with ${41.7}\,\text{dB}$, despite operating on $64\times$ fewer pixels. In other words, the effective capacity we observe at ${256}{\times}{256}\text{px}$ is comparable to what one would expect around ${20}{\times}{20}\text{px}$, confirming that the architecture fails to utilize the available resolution.
Because this setup does not require robustness to geometric or valuemetric transformations and we consider only gray images, we can use the ${32}{\times}{32}\text{px}$ model to demonstrate that higher capacities are possible. A simple tiling strategy suffices: each tile is embedded with an independent secret using the same model. The decoder similarly is applied per patch with the individual decoded messages concatenated to obtain the final combined message. Using ${256}{\times}{256}\text{px}$ as the reference size, tiling yields $64\times$ the capacity of the base ${32}{\times}{32}\text{px}$ model. Thus, tiling the 512-bit model—which already achieves 100% accuracy at ${41.7}\,\text{dB}$—produces a watermark with ${32{,}768}\,\text{bits}$ total capacity while maintaining the same PSNR (which is resolution-independent). This effective capacity of ${32{,}768}\,\text{bits}$ is already much closer to our bound of roughly ${600{,}000}\,\text{bits}$, though still only about ${0.167}\,\text{bpp}$.
It is interesting that the model could not learn at the ${256}{\times}{256}\text{px}$ resolution even for 1024 bits when it is clear that it is possible to embed 32,768 bits as seen here. More importantly, this shows that our bounds are not that far off and capacities of at least 32,768 bits are indeed possible.